OpenStudy (anonymous):

List the next three terms in the following sequence:

OpenStudy (wmj259):

Whats the sequence?

OpenStudy (anonymous):

1, 2, 4, 8, 16, 32,

OpenStudy (wmj259):

Notice the rate of which the terms change is *2 1*2=2 2*2=4 4*2=8 8*2=16 16*2=32 32*2=64 64*2=128 128*2=256

OpenStudy (anonymous):

thank you

OpenStudy (wmj259):

You are most certainly welcome. Please tell me if you understood the process?